If the median mass of 5 people in a lift is 65kg and a 70kg man enters then:

i) The new median will be about 66kg is it true and why


1
Expert's answer
2021-09-01T11:52:35-0400

At first we have mass of 5 people


"m_1\leq m_2\leq65\leq m_4\leq m_5"

Now a 70kg man enters

If "m_4\geq70," we have


"m_1\leq m_2\leq65<70\leq m_4\leq m_5"

"median_{new}=\dfrac{65+70}{2}=67.5"

If "65\leq m_4\leq70," we have


"m_1\leq m_2\leq65\leq m_4\leq 70\leq m_5"

"median_{new}=\dfrac{65+m_4}{2}"

"65\leq median_{new}\leq67.5"

Depending on value "m_4" the new median will be


"65\leq median_{new}\leq67.5"

I think that the statement is False.